OpenStudy (swag):

Y=1/2x+2 and Y=1/3x-4 How many solutions with this have and why

OpenStudy (swag):

@tafkas77

OpenStudy (anonymous):

Erm... this could have many solutions, depending on what it is you're studying right now. :) Is this a system of equations or...?

OpenStudy (amistre64):

difference slopes HAVE to have a single solution

OpenStudy (anonymous):

Well, there you go!^^ I was trying to figure out what exactly the problem was asking, but I guess amistre64 didn't need to, lol. I would go with what he said, @SWAG

OpenStudy (amistre64):

the "and" gave it away :)

OpenStudy (anonymous):

*facepalm* definitely did not pay attention to that

OpenStudy (swag):

x+2y=4 and 2x+3y=7 Solve by substitutiion how do i do this?

OpenStudy (anonymous):

@amistre64 Do you want to take this? I have to go pick up my little brother from school. @SWAG unless you want to wait :)

OpenStudy (anonymous):

Aaaaaand... he's gone. Let me see what I can do... you know what? I can drive and get my brother. I usually walk, but let me help you with this one.

OpenStudy (anonymous):

GAH I'm trying! Now let me type this out, kay? Sheesh lol

OpenStudy (anonymous):

wait, the problem is this? Y=1/2x+2 and Y=1/3x-4

OpenStudy (swag):

x+2y=4 and 2x+3y=7 Solve by substitutiion how do i do this?

OpenStudy (amistre64):

there is no time limits here :)

OpenStudy (anonymous):

Do you know where to start, or are you completely lost?

OpenStudy (anonymous):

@amistre64 :)

OpenStudy (amistre64):

my internet time is ending today so yall have fun

OpenStudy (anonymous):

He's not answering... and I need to get my little brother now; can you please help SWAG @Hero

OpenStudy (swag):

Never mind no rush

OpenStudy (swag):

Go ahead @tafkas77 it is ok

OpenStudy (swag):

:)

OpenStudy (anonymous):

Thanks. I'm sorry about that. :/

OpenStudy (swag):

It is ok :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@SWAG yes? :)

OpenStudy (swag):

could you answer now